# Network-Realized Model Predictive Control Part II: Distributed Constraint Management

### v1

@@ 7.1.1 The Dynamics of the Application
In order to showcase the effectiveness of the proposed design framework, we now focus on a numerical example centred around the vehicle platoon application discussed in [8]. We consider a formation of $N=10$ identical vehicles, in which the continuous-time models from (22) in [8] are configured with the following parameters: $m_{i}=1$, $\tau_{i}=0.1$ and $\sigma_{i}=1,\,\forall\,i\in\{1:N\}$. The resulting zero-order hold discretized dynamics of each car, for a sampling period of $T_{s}=0.1$, are given by
$\scriptsize\begin{bmatrix}p_{i}[k+1]\\ v_{i}[k+1]\\ \mu_{i}[k+1]\end{bmatrix}=\underbrace{\left[\scriptsize\begin{array}[]{rrr}1&0.1&-0.0331\\ 0&1&-0.5689\\ 0&0&0.3679\end{array}\right]}_{A_{car}}\scriptsize\begin{bmatrix}p_{i}[k]\\ v_{i}[k]\\ \mu_{i}[k]\end{bmatrix}+\underbrace{\scriptsize\begin{bmatrix}0.0381\\ 0.6689\\ 0.6321\\ \end{bmatrix}}_{B_{car}}u_{i}[k],\normalsize\vspace{-3mm}$
where $u_{i}$ is the control signal of the $i^{\text{th}}$ car, $p_{i}[k]$ denotes its (absolute) position, $v_{i}[k]$ is its speed and $\mu_{i}[k]$ is the state of the $i^{\text{th}}$ car’s actuator (the $\Phi$ terms in [8]).
@@
$\scriptsize\begin{bmatrix}p_{0}[k+1]\\ \ell_{v}[k+1]\end{bmatrix}=I_{N+1}\scriptsize\begin{bmatrix}p_{0}[k]\\ \ell_{v}[k]\end{bmatrix}+e_{1}\Delta p_{0}[k].\normalsize\vspace{-3mm}$
@@
7.1.2 First Layer Objectives and Platoon Model
In keeping with the control laws proposed in [8], we wish to designate each car as an individual area and to obtain a first layer in which the control signals are computed (in the absence of exogenous or second layer signals) as
$\mathrm{row}_{i}(u_{f}[k])=\left\{\tiny\begin{array}[]{l}\mathbf{\Gamma}_{1}(z)\star(y_{1}[k]+hv_{1}[k]),\ i=1,\\ \mathbf{\Gamma}_{i}(z)\star(y_{i}[k]\,+\,hv_{i}[k])+\mathbf{\Phi}_{i}(z)\star\mathrm{row}_{i-1}(u_{f}[k]),\end{array}\right.\normalsize$ (49)
for all $i\in\{2:N\}$, where $y_{i}[k]:=p_{i}[k]+\ell_{i}[k]-p_{i-1}[k]$ stands for the $i^{\text{th}}$ car’s length-based interdistance and $h>0$ is the platoon’s time headway (see [8] and the references therein for its theoretical justification), which has been chosen as $h=5$ for the considered application.
@@ 7.1.2 First Layer Objectives and Platoon Model
$T=\tiny\begin{bmatrix}I_{N+1}&O&O&O&\dots&O&O\\ T_{\ell 1}&I_{3}&O&O&\dots&O&O\\ T_{\ell 2}&T_{\ell d}&I_{3}&O&\dots&O&O\\ T_{\ell 3}&O&T_{\ell d}&I_{3}&\dots&O&O\\ \vdots&\vdots&&\ddots&\ddots&\vdots&\vdots\\ T_{\ell N}&O&O&O&\dots&T_{\ell d}&I_{3}\end{bmatrix}\in\mathbb{R}^{(4N+1)\times(4N+1)},\vspace{-3mm}$
@@
$T_{\ell i}=\left\{\scriptsize\begin{array}[]{l}\scriptsize\begin{bmatrix}e_{2}-e_{1}&O&O\end{bmatrix}^{\top}\in\mathbb{R}^{3\times(N+1)},\ i=1,\\ \vspace{-4mm}\\ \scriptsize\begin{bmatrix}e_{i+1}&O&O\end{bmatrix}^{\top}\in\mathbb{R}^{3\times(N+1)},\,\forall\,i\in\{2:N\}.\end{array}\right.\vspace{-3mm}$
to obtain the platoon’s model of type (4a), given by
$x[k+1]=Ax[k]+Bu[k]+e_{1}\Delta p_{0}[k],\vspace{-3mm}$ (50)
@@ 5. $e)$
    $B_{d}:=e_{1}\in\mathbb{R}^{4N+1}$ and $d[k]:=\Delta p_{0}[k]$, by direct correspondence with (4a).
@@
[…]
By choosing $\mathbb{C}_{g}=\{z\in\mathbb{C}\ |\ |z|<1\}$, the aim of the first layer will be to stabilize platoon’s closed-loop dynamics, in the sense that, by implementing the control laws of type (49) as detailed in our companion paper [10], the closed-loop dynamics will be described by
$\scriptsize\begin{bmatrix}p_{0}[k+1]\\ \ell_{v}[k+1]\\ x_{f}[k+1]\end{bmatrix}=\scriptsize\begin{bmatrix}1&O&O\\ O&I_{N}&O\\ O&O&A_{f}\end{bmatrix}\scriptsize\begin{bmatrix}p_{0}[k]\\ \ell_{v}[k]\\ x_{f}[k]\end{bmatrix}+\scriptsize\begin{bmatrix}B_{0}\\ B_{\ell}\\ B_{f}\end{bmatrix}d_{s}[k],\normalsize\vspace{-3mm}$ (51)
@@ 2. $b)$
    $x_{f}=\begin{bmatrix}x_{1}^{\top}[k]&w_{1}^{\top}[k]&\dots&x_{N}^{\top}[k]&w_{N}^{\top}[k]\end{bmatrix}^{\top}$, with $w_{i}[k]$ being the state vectors belonging to the NRF implementations of type (8), for all $i\in\{1:N\}$;
@@ 3. $c)$
    $d_{s}$ is the signal vector from (13).
@@
Since $p_{0}[k]$ and $\ell_{i}[k]$, for any $i\in\{1:N\}$, cannot be influenced by the $u_{i}[k]$ signals, the first layer will be unable to stabilize their dynamics. However, given the nature of our application and the physical interpretation of the variables, the aforementioned stabilization is of no practical interest. The second layer will focus only upon the identity $x_{f}[k+1]=A_{f}x_{f}[k]+B_{f}d_{s}[k]$, which will represent our first layer’s closed-loop dynamics.
7.1.3 Second Layer Objectives
The primary objective of the second layer is to ensure constraint satisfaction, with respect to the variables that make up $x_{i}[k]$ and $u_{i}[k]$, in the presence of recursive feasibility. Thus, for all $i\in\{1:N\}$ and all $k\geq k_{0}$, we wish to have that:
$y_{i}[k]\in[-360,0],\ v_{i}[k]\in[0,36],\ u_{i}[k]\in[-10,10].\vspace{-3mm}$ (52)
These constrains must be satisfies while employing the same communication scheme as the one considered in (49), i.e., the $i^{\text{th}}$ car can employ only information that is accessed locally or from the car directly in front of it.
Additionally, due to the fact that the prototype NRF controller from [8] (if properly tuned) ensures a high level of performance, we want the second layer to intervene as little as possible in the closed-loop dynamics of the first one. In other words, most of the control action should be undertaken by the $u_{fi}[k]$ signals, whereas the entries of $u_{s1i}[k]$ and $u_{s2i}[k]$ should have negligible absolute values (compared to those of $u_{fi}[k]$) whenever the constraints from (52) are not in danger of being violated.
@@ 7.2.1 The First Layer
Since the application considered in this numerical example already has a dedicated NRF-based solution, we defer to [8] for the procedural details regarding the design of our first layer. In so doing, we briefly discuss the numerical implementation of these control laws, before focusing our attention on the design of the second layer.
For the homogeneous platoon whose physical parameters were given in the previous subsection, the first layer command signals will be computed as stated in (11), via the state-space implementations of type (8) given by
@@
$u_{fi}[k]=\left[\scriptsize\begin{array}[]{c|cc}a_{i}-z&b_{\mathbf{\Phi}i}&B_{\mathbf{\Gamma}i}\\ \hline\cr 1&0&O\end{array}\right]\star\tiny\begin{bmatrix}u_{f(i-1)}[k]+\beta_{f(i-1)}[k]\\ x_{i}[k]+\zeta_{i}[k]+u_{s1i}[k]+\beta_{s1i}[k]\end{bmatrix},\normalsize$ (54)
@@
Table 1: Coefficients of the first layer controller implementations
By using the state-space systems in (50) and (53)-(54) to form the first layer’s closed-loop system, we obtain a realization of type (51) in which the spectral radius of $A_{f}$ is $0.9936$, thus validating all of the design objectives associated with the first layer. In addition to this, the resulting closed-loop dynamics of each car in the platoon will be given by the realizations from (55a)-(55b), which are located at the top of this page and in which
$A_{w}=\left[\scriptsize\begin{array}[]{ccc}0&\ -0.1&\ 0.0331\\ 0&\ 0&\ 0\\ 0&\ 0&\ 0\\ \end{array}\right]\text{ and }B_{w}=\left[\scriptsize\begin{array}[]{c}-0.0381\\ 0\\ 0\\ \end{array}\right].\vspace{-3mm}$
@@ 7.2.2 The Second Layer
The first step in designing our control architecture’s second layer, with the aim of ensuring its recursive feasibility, consists in defining all of the sets that were discussed in Section 5. We begin by addressing the command signals generated by the MPC layer, for which we impose:
$\scriptsize\left\{\begin{array}[]{l}\mathcal{U}_{s1i}=\left\{u_{s1i}\in\mathbb{R}^{3}\,\big{|}\,\left|e^{\top}_{1}u_{s1i}\right|\leq 720,\ \left|e^{\top}_{2}u_{s1i}\right|\leq 72,\right.\\ \left.\qquad\qquad\qquad\qquad\qquad\qquad\qquad e^{\top}_{3}u_{s1i}=0\right\},\,\forall\,i\in\{1:N\},\\ \mathcal{U}_{s2i}=\left\{u_{s2i}\in\mathbb{R}\,|\,\left|u_{s2i}\right|\leq 5\right\},\,\forall\,i\in\{1:N\}.\end{array}\right.\vspace{-3mm}$
@@
and, finally, we address the constraints related to the platoon’s states by first introducing the matrix denoted $W_{x}:=\mathrm{row}_{1}\left(\tiny\begin{bmatrix}A_{car}&B_{car}\end{bmatrix}\right)\mathrm{diag}(0,I_{3})$ and then constructing the set $\widehat{\mathcal{X}}\subseteq\mathbb{R}^{4}$ given in (56), which is located at the top of this page. By choosing $T_{i}=1$ for all $i\in\{1:N\}$ and by employing $\widehat{\mathcal{X}}$, we adapt Algorithm 10.5 from [4] to the dynamics from (55a)-(55b), while also accounting for the disturbance discussed in Section 4.2.
@@
The sets which are obtained through the aforementioned adaptation are the robust control invariant sets $\,\mathcal{C}_{i}\subseteq\widehat{\mathcal{X}}$, for all $i\in\{1:N\}$. Due to our particular choice of $T_{i}=1$, these sets [^2] The numerical procedures used to compute $\mathcal{C}_{i}$, along with all of the data sets discussed in this numerical example, are available at https://github.com/AndreiSperila/NRMPC act as low-complexity alternatives to the ones from (43), while also satisfying (by their very construction, see Section 10.4 in [4]) a condition which is a direct analogue to (48), thus guaranteeing the recursive feasibility of our second layer.
In light of the arguments made in Remark 7.2, we proceed to replace the constraints $\xi_{i}[k_{0}+1]\in\Xi_{i1}$ in (42) with $\scriptsize\begin{bmatrix}x_{i}^{\top}[k_{0}+1]&w_{i}[k_{0}+1]\end{bmatrix}^{\top}\in\mathcal{C}_{i},\,\forall\,i\in\{1:N\}$. We then choose $\overline{T}_{i}=0$ and we configure the costs in (42) as
$\scriptsize\begin{array}[]{l}10^{-9}y_{i}^{2}[k_{0}+1]+u_{s2i}^{2}[k_{0}]+u_{s1i}^{\top}[k_{0}]u_{s1i}[k_{0}]\end{array}\normalsize,\vspace{-3mm}$ (57)
thus concluding the design phase of the second layer.
@@ 7.3 Simulation Scenario and Results
[…]
is the speed of the virtual $0^{\text{th}}$ car and $\mathds{1}[k]$ denotes the discrete-time Heaviside step function. As can be seen in Figure 3 below, in which $v_{0}[k]$ is depicted in orange, the speed of each car in the platoon will try to match that of the $0^{th}$ car in order to achieve steady-state.
@@
With respect to the limitations placed upon the signals being sent to the car actuators, our control scheme is once again successful. As can be seen from Figures 5 and 6 below, the sum of the signals from the two aforementioned figures is well within the bounds specified in (52), even taking into account the amplitudes of the entries belonging to $\beta_{s2}[k]$
Figure 5: Outputs of the first layer subcontrollers
Figure 6: Second layer outputs to the platoon
@@
Figure 7: Norm values for second layer outputs to the first layer subcontrollers
@@
The brief intervals in which the two signal vectors are of the same order of magnitude as the first layer control signals are those in which the second layer intervenes in order to saturate either the output of the first car’s NRF-based subcontroller (see Figure 8 below), or the speed of the lead car in the platoon (see Figure 9 at the top of the next page’s left-hand column) with respect to the bounds prescribed to these variables.
@@
Figure 8: Constraint satisfaction for the first layer command of the platoon’s first car (second layer subcontroller active in the green shaded area)
@@
Figure 9: Constraint satisfaction for the speed of the platoon’s first car (second layer subcontroller active in the green shaded area)

### v4

@@ 7.1.1 The Dynamics of the Application
In order to showcase the effectiveness of the proposed design framework, we now focus on a numerical example centred around the vehicle platoon application discussed in [plutonizare]. We consider a formation of $N=10$ identical vehicles, in which the continuous-time models from (22) in [plutonizare] are configured with the following parameters: $m_{i}=1$, $\tau_{i}=0.1$ and $\sigma_{i}=1,\,\forall\,i\in\{1:N\}$. The resulting zero-order hold discretised dynamics of each car, for a sampling period of $T_{s}=0.1$, are given by
$\scriptsize\begin{bmatrix}p_{i}[k+1]\\ v_{i}[k+1]\\ \mu_{i}[k+1]\end{bmatrix}=\underbrace{\left[\scriptsize\begin{array}[]{rrr}1&0.1&-0.0331\\ 0&1&-0.5689\\ 0&0&0.3679\end{array}\right]}_{A_{car}}\scriptsize\begin{bmatrix}p_{i}[k]\\ v_{i}[k]\\ \mu_{i}[k]\end{bmatrix}+\underbrace{\scriptsize\begin{bmatrix}0.0381\\ 0.6689\\ 0.6321\\ \end{bmatrix}}_{B_{car}}u_{i}[k],\normalsize\vskip-8.53581pt$
where $u_{i}$ is the control signal of the $i^{\text{th}}$ car, $p_{i}[k]$ denotes its (absolute) position, $v_{i}[k]$ is its speed and $\mu_{i}[k]$ is the state of the $i^{\text{th}}$ car’s actuator (the $\Phi$ terms in [plutonizare]).
@@
$\scriptsize\begin{bmatrix}p_{0}[k+1]\\ \ell_{v}[k+1]\end{bmatrix}=I_{N+1}\scriptsize\begin{bmatrix}p_{0}[k]\\ \ell_{v}[k]\end{bmatrix}+e_{1}\Delta p_{0}[k].\normalsize\vskip-8.53581pt$
@@
7.1.2 First-Layer Objectives and Platoon Model
In keeping with the control laws proposed in [plutonizare], we wish to designate each car as an individual area and to obtain a first layer in which the control signals are computed (in the absence of exogenous or second-layer signals) as
[…]
for all $i\in\{2:N\}$, where $y_{i}[k]:=p_{i}[k]+\ell_{i}[k]-p_{i-1}[k]$ stands for the $i^{\text{th}}$ car’s length-based interdistance and $h>0$ is the platoon’s time headway (see [plutonizare] and the references therein for its theoretical justification), which has been chosen as $h=5$ for the considered application.
@@ 7.1.2 First-Layer Objectives and Platoon Model
$T=\tiny\begin{bmatrix}I_{N+1}&O&O&O&\dots&O&O\\ T_{\ell 1}&I_{3}&O&O&\dots&O&O\\ T_{\ell 2}&T_{\ell d}&I_{3}&O&\dots&O&O\\ T_{\ell 3}&O&T_{\ell d}&I_{3}&\dots&O&O\\ \vdots&\vdots&&\ddots&\ddots&\vdots&\vdots\\ T_{\ell N}&O&O&O&\dots&T_{\ell d}&I_{3}\end{bmatrix}\in\mathbb{R}^{(4N+1)\times(4N+1)},\vskip-8.53581pt$
@@
[…]
to obtain the platoon’s model of type (1a)-(1c), given by
$x[k+1]=Ax[k]+Bu[k]+e_{1}\Delta p_{0}[k],\vskip-8.53581pt$ (48)
@@
    $B_{d}:=e_{1}\in\mathbb{R}^{4N+1}$ and $d[k]:=\Delta p_{0}[k]$, by direct correspondence with (1a)-(1c).
@@
$\scriptsize\begin{array}[]{l}\widehat{\mathcal{X}}:=\left\{\widehat{x}\in\mathbb{R}^{4}\,\big|\,-360\leq e_{1}^{\top}\widehat{x}\leq 0,\ 0\leq e_{2}^{\top}\widehat{x}\leq 36,\ \left|e_{3}^{\top}\widehat{x}\right|\leq 10,\ \left|e_{4}^{\top}\widehat{x}\right|\leq 4.99,\ 5.01\left|e_{1}^{\top}B_{car}\right|\leq W_{x}\widehat{x}\leq 36T_{s}-5.01\left|e_{1}^{\top}B_{car}\right|\right\}.\end{array}\normalsize\vskip-11.38109pt$ (54)
By choosing $\mathbb{C}_{g}=\{z\in\mathbb{C}\ |\ |z|<1\}$, the first layer will aim to stabilise platoon’s closed-loop dynamics, in the sense that, by implementing the control laws of type (47) as detailed in our companion paper [part1], the closed-loop dynamics will be described by
$\scriptsize\begin{bmatrix}p_{0}[k+1]\\ \ell_{v}[k+1]\\ x_{f}[k+1]\end{bmatrix}=\scriptsize\begin{bmatrix}1&O&O\\ O&I_{N}&O\\ O&O&A_{f}\end{bmatrix}\scriptsize\begin{bmatrix}p_{0}[k]\\ \ell_{v}[k]\\ x_{f}[k]\end{bmatrix}+\scriptsize\begin{bmatrix}B_{0}\\ B_{\ell}\\ B_{f}\end{bmatrix}d_{s}[k],\normalsize\vskip-8.53581pt$ (49)
@@
    $x_{f}[k]=\begin{bmatrix}x_{1}^{\top}[k]&w_{1}^{\top}[k]&\dots&x_{N}^{\top}[k]&w_{N}^{\top}[k]\end{bmatrix}^{\top}$, with $w_{i}[k]$ being the state vectors belonging to the NRF implementations of type (10), for all $i\in\{1:N\}$;
@@
    $d_{s}[k]$ is the signal vector from (7).
@@
Since $p_{0}[k]$ and $\ell_{i}[k]$, for any $i\in\{1:N\}$, cannot be influenced by the $u_{i}[k]$ signals, the first layer will be unable to stabilise their dynamics. However, given the nature of our application and the physical interpretation of the variables, the aforementioned stabilisation is of no practical interest. The second layer will focus only on the identity $x_{f}[k+1]=A_{f}x_{f}[k]+B_{f}d_{s}[k]$, which will represent our first layer’s closed-loop dynamics.
7.1.3 Second-Layer Objectives
The main goal of the second layer is to ensure constraint satisfaction, with respect to the variables that make up $x_{i}[k]$ and $u_{i}[k]$, in the presence of recursive feasibility. Thus, for all $i\in\{1:N\}$ and all $k\geq k_{0}$, we wish to have:
$y_{i}[k]\in[-360,0],\ v_{i}[k]\in[0,36],\ u_{i}[k]\in[-10,10].\vskip-8.53581pt$ (50)
These constraints must be satisfied while employing the same communication scheme as the one considered in (47), i.e., the $i^{\text{th}}$ car can employ only information that is accessed locally or from the car directly in front of it.
Additionally, due to the fact that the prototype NRF controller from [plutonizare] (if properly tuned) ensures a high level of performance, we want the second layer to intervene as little as possible in the closed-loop dynamics of the first one. In other words, most of the control action should be undertaken by the $u_{fi}[k]$ signals, whereas the entries of $u_{s1i}[k]$ and $u_{s2i}[k]$ should have negligible absolute values (compared to those of $u_{fi}[k]$) whenever the constraints from (50) are not in danger of being violated.
@@ 7.2.1 The First Layer
Since the application considered in this numerical example already has a dedicated NRF-based solution, we defer to [plutonizare] for the procedural details regarding the design of our first layer. In so doing, we briefly discuss the numerical implementation of these control laws, before focusing our attention on the design of the second layer.
For the homogeneous platoon whose physical parameters were given in the previous subsection, the first-layer command signals will be computed as stated in (12), via the state-space implementations of type (10) given by
@@
[…]
@@
Table 1: Coefficients of the first-layer controller implementations
By using the state-space systems in (48) and (51)-(52) to form the first layer’s closed-loop system, we obtain a realisation of type (49) in which the spectral radius of $A_{f}$ is $0.9936$, thus validating all of the design objectives associated with the first layer. In addition to this, the resulting closed-loop dynamics of each car in the platoon will be given by the realisations from (53a)-(53b), which are located at the top of this page and in which
$A_{w}=\left[\scriptsize\begin{array}[]{ccc}0&\ -0.1&\ 0.0331\\ 0&\ 0&\ 0\\ 0&\ 0&\ 0\\ \end{array}\right]\text{ and }B_{w}=\left[\scriptsize\begin{array}[]{c}-0.0381\\ 0\\ 0\\ \end{array}\right].\vskip-8.53581pt$
@@ 7.2.2 The Second Layer
The first step in designing our control architecture’s second layer, with the aim of ensuring its recursive feasibility, consists of defining all of the sets discussed in Section 5. We begin by addressing the command signals
generated by the MPC layer, for which we impose:
[…]
@@
and, finally, we address the constraints related to the platoon’s states by first introducing the matrix denoted $W_{x}:=\mathrm{row}_{1}\left(\tiny\begin{bmatrix}A_{car}&B_{car}\end{bmatrix}\right)\mathrm{diag}(0,I_{3})$ and then constructing the set $\widehat{\mathcal{X}}\subseteq\mathbb{R}^{4}$ given explicitly in (54), which is located at the top of this page.
car number 1 2 3 4 5 6 7 8 9 10
time (seconds) 211.740 757.767 390.860 396.871 295.800 235.603 155.952 204.719 74.838 71.048
Table 2: Computation times associated with second-layer synthesis
By choosing $T_{i}=1$ for all $i\in\{1:N\}$ and by leveraging $\widehat{\mathcal{X}}$, while also accounting for the disturbance discussed in Section 4.3, we proceed to implement our design procedure. The resulting numerical routines and all of the data sets from this numerical example are available at https://github.com/AndreiSperila/NRMPC. Moreover, Table 2 (located at the top of the next page) showcases the computation times associated with the complete design of every car’s second-layer subcontroller.
@@
As explained in the comments that accompany the source code from the above-indicated link, our approach entails the implicit design of the robust control invariant sets $\mathcal{C}_{i}\subseteq\widehat{\mathcal{X}}$, which (by their very construction; see, for example, Section 10.4 in [BBM]) guarantee the recursive feasibility of our second layer. We stress the fact that (as previously indicated in Section 6) the sets $\mathcal{C}_{i}$ are not computed explicitly, with their existence being a desirable and natural consequence of our proposed approach.
Finally, we choose $\overline{T}_{i}=0$ and we set the costs in (8) as
$\begin{array}[]{l}10^{-9}y_{i}^{2}[k+1]+u_{s2i}^{2}[k]+u_{s1i}^{\top}[k]u_{s1i}[k]\end{array}\normalsize,\vskip-8.53581pt$ (55)
thereby concluding the design phase of the second layer.
@@ 7.3 Simulation Scenario and Results
$\scriptsize\begin{array}[]{l}v_{0}[k]=10\cdot\mathds{1}[k]\,\textbf{--}\,7\cdot\mathds{1}[k\,\textbf{--}\,400]+30\cdot\mathds{1}[k\,\textbf{--}\,1200]\,\textbf{--}\,30\cdot\mathds{1}[k\,\textbf{--}\,1300]\end{array}\normalsize\vskip-8.53581pt$
is the speed of the virtual $0^{\text{th}}$ car and $\mathds{1}[k]$ denotes the discrete-time Heaviside step function. As can be seen in Figure 3 below, in which $v_{0}[k]$ is depicted in orange, the speed of each car in the platoon will try to match that of the $0^{th}$ car in order to achieve steady-state evolution.
@@
With respect to the limitations placed upon the signals sent to the car actuators, our control scheme is once again successful. As can be seen from Figures 5 and 6 below, the sum of the signals from the two aforementioned figures is well within the bounds from (50), even taking into account the amplitudes of the entries in $\beta_{s2}[k]$.
Figure 5: Outputs of the first-layer subcontrollers
Figure 6: Second-layer outputs to the platoon
@@
Figure 7: Norm values for second-layer outputs to the first-layer subcontrollers
@@
The brief intervals in which the two signal vectors are of the same order of magnitude as the first-layer control signals are those in which the second layer intervenes in order to saturate either the output of the first car’s NRF-based subcontroller (see Figure 8 below), or the speed of the lead car in the platoon (see Figure 9, also located below) with respect to these variables’ prescribed bounds.
@@
Figure 8: Constraint satisfaction for the first-layer command of the platoon’s first car (second-layer subcontroller active in the green shaded area)
@@
Figure 9: Constraint satisfaction for the speed of the platoon’s first car (second-layer subcontroller active in the green shaded area)
Finally, as anticipated in Section 6.3.3 and as showcased by the runtime statistics showcased in Tables 3 and 4 above, the implementation costs associated with our proposed solution are sufficiently modest such that command computation is well below the sampling period of our control system, namely $T_{s}=0.1$ seconds. Note that all simulations associated with these runtimes were performed via MATLAB R2024b, on a mid-range laptop equipped with an i9-100885H Intel Core processor.
car number min max mean median mode
1 0.003 0.213 0.005 0.004 0.004
2 0.001 0.057 0.003 0.002 0.002
3 0.001 0.023 0.002 0.002 0.002
4 0.001 0.022 0.002 0.002 0.002
5 0.001 0.022 0.002 0.002 0.002
6 0.001 0.233 0.002 0.002 0.002
7 0.001 0.008 0.002 0.002 0.002
8 0.001 0.010 0.002 0.002 0.002
9 0.001 0.013 0.002 0.002 0.002
10 0.001 0.011 0.002 0.002 0.002
Table 3: NRF-based subcontroller runtime statistics in milliseconds
car number min max mean median mode
1 1.678 4.375 2.202 2.127 1.872
2 1.688 5.028 2.028 1.955 1.877
3 1.659 4.347 1.949 1.884 1.881
4 1.661 3.882 1.930 1.868 1.742
5 1.648 3.929 1.921 1.861 1.853
6 1.630 4.936 1.931 1.864 1.826
7 1.629 4.895 1.913 1.848 1.807
8 1.624 4.829 1.924 1.852 1.852
9 1.639 4.233 1.956 1.838 1.847
10 1.614 4.680 1.989 1.838 1.786
Table 4: MPC-based subcontroller runtime statistics in milliseconds
7.4 Comparison with Alternative Solutions
In order to conclude this numerical example, we present a comparison with several prominent alternative techniques from literature, which tackle the platooning problem in an MPC-based framework. We refer here to the procedures proposed in [plat_MPC_1, plat_MPC_2, plat_MPC_3, plat_MPC_4, plat_MPC_5] and we point out that a major limitation for all of them is the fact that, unlike our proposed solution, no source of disturbance (either from exogenous sources or from model uncertainty) is considered in any of these approaches. Moreover, the control laws presented in [plat_MPC_1, plat_MPC_2, plat_MPC_3] all share in common the fact that they do not explicitly support constraints for the state variables of the platoon’s cars, except for a restrictive use of terminal constraints, which are meant to ensure a number of theoretical guarantees.
Another common feature of the cited works is the fact that the lead cars in [plat_MPC_2, plat_MPC_3, plat_MPC_4] must broadcast information to all other vehicles in the formation, which places an impractical amount of strain on the communication infrastructure of the platoon. In addition to this fact, the property of asymptotic closed-loop stability is retrieved in [plat_MPC_1, plat_MPC_3, plat_MPC_4] by carefully tuning the stage costs of the MPC-based subcontrollers, rather than it being an inbuilt feature, as ensured by the NRF-based first layer (see [part1] and [plutonizare]). This being said, we now address the specific aspects of the five aforementioned techniques:
  1. a)
    One of the key features of the procedure from [plat_MPC_1] is the fact that the above-mentioned terminal constraint requires the computation of a regional average value, evaluated with respect to a number of neighbouring cars. In addition to the demands this places on the communication infrastructure of the platoon, the feasibility of the resulting control problem is noticeably dependent on choosing a large enough prediction horizon. While the overall computational complexity of the optimisation problem to be solved is similar to that of a QP-based formulation in (8), increasing the prediction horizon may have a significant computational impact which can be bypassed in our framework (recall Remark 6.6).
  2. b)
    The main limitation of the technique in [plat_MPC_2] is the fact that the optimisation problem associated with the MPC-based subcontrollers has to be run twice during a car’s command computation routine, with said problem having a comparable cost to a QP-based formulation of our problem in (8). Furthermore, we point out that both the asymptotic and string stability of the closed-loop system are only guaranteed when the platoon’s initial condition satisfies a series of restrictive assumptions.
  3. c)
    In addition to the list of common restrictions detailed above, the solution proposed in [plat_MPC_3] is affected by another limitation; namely, that it can only be initialised when the platoon is already in equilibrium. In spite of these disadvantages, we point out that the procedure’s computational cost is appealingly modest, being roughly analogous to that of our own procedure’s QP-based formulation, since it does not require the double iteration from [plat_MPC_2].
  4. d)
    The approach from [plat_MPC_4] possesses some structural similarity with the one presented in this paper, in the sense that it includes a rudimentary, PID-based lower control layer, while also enabling the use of limited, box-based constraints for the platoon’s states. However, we point out that the lower-level PIDs are not only unable to ensure closed-loop string stability (in contrast to our first layer), but the NRF-based subcontrollers offer many more degrees of freedom, in terms of both sparsity patterns and closed-loop performance (see [part1]). Since [plat_MPC_4] is heavily inspired by the approach presented in [plat_MPC_3], the same computational benefits apply to it as well.
  5. e)
    Finally, the procedure laid out in [plat_MPC_5] proposes a centralised take on the problem and, unsurprisingly, it showcases a high degree of closed-loop performance in the presence of state constraints. In addition to being intractable in the distributed setting, its critical aspect is given by the fact that closed-loop string stability is incompatible with the aforementioned constraints. In contrast to our solution, when safety-critical constraints become active, the technique given in [plat_MPC_5] sacrifices the platoon’s string stability for guaranteed constraint satisfaction.